Question 1

  1. Unless otherwise stated, assume all times reported refer to averages from exponential distributions and that we are looking at stable processes.

If the average time between arrivals is 10 minutes, what is the arrival rate?

a.

6 jobs per hour

b.

0.1 jobs per minute

c.

0.001666 jobs per second

d.

All of the above

1 points

Question 2

  1. For a system with a single server, if the arrival rate is six jobs per hour and the average service time is 5 minutes, what is utilization?

a.

5/6

b.

5/10

c.

6 jobs / 5 minutes

d.

Cannot be found

1 points

Question 3

  1. For a system with two servers, if the arrival rate is six jobs per hour and the average service time for each server is 5 minutes what is utilization?

a.

5/6

b.

5/10

c.

6/(6 * 2)

d.

(6 * 2)/6

1 points

Question 4

  1. For a system with two servers, if the arrival rate is 12 jobs per hour and the average service rate for each server is eight jobs per hour, what is utilization?

a.

8/12

b.

(8*2)/12

c.

12/8

d.

12 / (8*2)

1 points

Question 5

  1. Little's Law teaches that on average (# of jobs in the system. = Throughput *Cycle time. If L is the length of the line including the customer being served, and W is the time in the system including the service time and we have a throughput rate of Lambda, then what does Little's Law say about this system?

a.

Lambda = L * W

b.

L * Lambda = W

c.

W = L * Lambda

d.

L = Lambda * W

1 points

Question 6

  1. What is the general relationship between utilization ( ) and queue length (L)? (Assume no change in the service rate.)

a.

As rises, L rises

b.

As rises L falls

c.

and L are independent

d.

It depends on utilization

1 points

Question 7

  1. As the COV of arrival times (Ci. increases what happens to average queue length (L)?

a.

As Ci rises L rises

b.

As Ci rises L falls

c.

Ci and L are independent

d.

It depends on utilization

1 points

Question 8

  1. Since the mean and variance of an exponential distribution are the same, what is the coefficient of variation for an exponential distribution?

a.

0

b.

1/2

c.

1

d.

I don't know

1 points

Question 9

  1. What happens to the average length of the line "L" if we move from a system that can accommodate an infinite queue length to one with a maximum queue length?

a.

It goes up

b.

It stays the same

c.

It goes down

d.

It depends on how large the variance is

1 points

Question 10

  1. If I only accommodate a finite number of waiting customers, what happens if I reduce the size of this buffer?

a.

More customers will be served

b.

More customers will be blocked

c.

More customers will be served, and more will be blocked

d.

Fewer customers will be blocked
Answer #1

As per policy only first question will be answered

1. Inter-arrival rate: 10 minutes

Hence per hour = 60/10 = 6 Therefore (a) is correct

Per minute = 6/60 = 0.1 per minute . Therefore option (b) is correct

Per second = per minute / 60 = 0.1/60 = 0.0016 , There option (c)

Hence option (d) is true